solve for x: sin(2x-3) csc(5x-15)=1
Answer by CharlesG2(828)

solve for x: sin(2x-3) csc(5x-15)=1

csc = 1/sin
sin(2x - 3)csc(5x - 15) = 1
sin(2x - 3)/sin(5x - 15) = 1
sin(2x - 3) = sin(5x - 15)
2x - 3 must equal 5x - 15
2x - 3 = 5x - 15
-3x = -12
x = -12/-3 = 12/3 = 4